# The genealogy of nearly critical branching processes in varying environment

### v3

@@ 4.3 The reduced process of a CPP
Recall the definition of the reduced process $(L_{s},0<s<t)$ of an ultrametric tree from Section 3.2.1. The process $L_{s}$ counts the number of ancestral lineages at time $s$, starting from the leaves at $s=0$ and going toward the root at $s=t$. We give a characterization of the reduced process associated to the environmental CPP. It is usual to express the reduced process of a branching process as a time-changed pure birth process, also known as a Yule process.
@@
Let $(L_{s},0<s\leq t)$ be the reduced process associated to the environmental CPP $[(0,Z_{e}),d_{\nu_{e}},\operatorname{Leb}]$. Then
@@
Consider a Brownian CPP, that is the random ultrametric space with distribution $\operatorname{CPP}_{1}(\tfrac{dx}{x^{2}},1)$, denoted by $[(0,Z_{B}),d_{B},\operatorname{Leb}]$, with associated reduced process $(L^{\prime}_{s},s\leq 1)$. It is well known that
@@
where $(Y(u),u\geq 0)$ is a Yule process, see for instance [10, Lemma 5.5] or [47, Theorem 2.3]. By Proposition 4.2, $[(0,Z_{B}),F^{-1}(d_{B}),\bar{\rho}_{t}\operatorname{Leb}]$ is distributed as the environmental CPP (30). Since the identity $F^{-1}(d(x,y))<t\iff d(x,y)<F(t)$ holds, the reduced process $(L_{s},0<s\leq t)$ of $[(0,Z_{B}),F^{-1}(d_{B}),\bar{\rho}_{t}\operatorname{Leb}]$ is
@@
and since $L_{t-s}=L^{\prime}_{F(t-s)}\stackrel{{\scriptstyle d}}{{=}}Y(-\log F(t-s))$ the result follows by noting that
$\forall s\leq t,\quad\frac{1}{F(t-s)}=\frac{2\bar{\rho}_{t}}{\int_{[s,t]}e^{X_{t}-X_{u}}\sigma^{2}(du)}.$
@@ 5.2 The many-to-few formula
where $d_{u}$ denotes the out-degree of $u$. An ultrametric tree with $k$ leaves can always be encoded as a sequence of $k-1$ elements of $\{1,\dots,n\}$ giving the depth of the successive coalescence times between the leaves. More precisely, let $\ell_{1},\dots,\ell_{k}$ be the leaves of $\tau$, ordered such that
@@
where $<$ is the planar (lexicographical) order of the tree. Then $\tau$ can be encoded as the sequence
@@
where $u\wedge v$ denotes the most recent common ancestor of $u$ and $v$. Conversely, given a sequence $(h_{1},\dots,h_{k-1})\in\{1,\dots,n\}$ one can find an ultrametric tree $\tau$ such that $\mathcal{H}(\tau)=(h_{1},\dots,h_{k-1})$, that is, $\mathcal{H}$ is a bijection. The tree $\mathcal{H}^{-1}(h_{1},\dots,h_{k-1})$ is obtained through the discrete CPP construction illustrated in Figure 3.
@@
Note that there are at most $k-1$ such branch points, and that we have the identity
@@
We now construct the $k$-spine tree using i.i.d. coalescence depths in the CPP construction. More precisely, for each $N$ consider a probability distribution $(p_{i},i\leq tN)$ on $\{1,\dots,tN\}$. (As in the rest of the paper we make the dependency on $N$ implicit). Let $(H^{(N)}_{i};\,i\geq 1)$ be a sequence of i.i.d. random variables with distribution $(p_{i},i\leq tN)$. The $k$-spine tree $S_{k}$ is defined as the random ultrametric tree $S_{k}=\mathcal{H}^{-1}(H^{(N)}_{1},\dots,H^{(N)}_{k-1})$. As in (18) introduce
@@
Following the notation in [15] we will denote the distribution of this array by $\mathbb{Q}^{k,N}$. Although $\mathbb{Q}^{k,N}$ is formally defined as a measure on $\{0,\dots,tN\}^{k\times k}$, by construction it corresponds to the matrix of pairwise distances between the leaves of the tree $S_{k}$. By abuse of notation, when convenient, we also think of $\mathbb{Q}^{k,N}$ the law of $S_{k}$, that is as a measure on ultrametric trees.
@@
The reader might be confused by the fact that $\mathbb{Q}$ denotes the law of the size-biased BPVE tree, whereas $\mathbb{Q}^{k,N}$ is the law of a tree $S_{k}$ with a fixed number $k$ of leaves. One could derive an extension of Lemma 5.1 that connects the law $\mathbb{P}$ (biased by its $k$-th factorial moment) to that of a tree with $k$ distinguished lineages obtained by grafting independent subtrees on the spine tree $S_{k}$, see [15, Theorem 3]. In [15] the distribution of this larger tree is denoted by ¯$\bar{\mathbb{Q}}^{k,N}$, and $\mathbb{Q}^{k,N}$ corresponds to the law of the tree spanned by the $k$ distinguished lineages. We follow this notation for easier comparison with [15], even though we do not need such a precise result here.
@@
where $f^{(d_{u})}_{|u|+1}$ denotes the $d_{u}$-th derivative of the generating function of the offspring distribution in generation $|u|$. Note that we have included the $k!$ term in the definition of $\Delta_{k}$ compared to [15]. (As in the rest of the paper, we make the dependence of $\Delta_{k}$ and $(p_{i},i\leq tN)$ on $N$ implicit, but keep it explicit for $\mathbb{Q}^{k,N}$ and $H^{(N)}_{i}$ for later purpose.)
@@ 6 Proof of the Kolmogorov estimate
We aim to prove that $q_{k}\to 1$ with a uniform bound for all $k\leq(1-\varepsilon)Nt$, which implies that it is sufficient to prove that the r.h.s. in (41) diverges uniformly to $\infty$ for all $k\leq(1-\varepsilon)Nt$. We have the estimate
@@
by Lemma 1 in [32]. The r.h.s above can be bounded from below via
$\sum_{j=(1-\varepsilon)Nt}^{Nt-1}\varphi_{j+1}(0)\geq c\sqrt{\kappa_{N}}(\sigma^{2}(t)-\sigma^{2}(t-\varepsilon)),$ (44)
by Lemma A.5. Therefore, recalling that by (6) $\mu_{j}$ is uniformly bounded, it follows for all $k\leq(1-\varepsilon)Nt$ that
$\frac{1}{1-q_{k}}=\frac{1}{\mathbb{P}\left(Z_{Nt}>0\mid Z_{k}=1\right)}\geq\sum_{j=(1-\varepsilon)Nt}^{Nt-1}\frac{\mu_{k}\varphi_{j}(0)}{\mu_{j}}\geq c^{\prime}\sqrt{\kappa_{N}}$ (45)
@@
$q_{k}\geq 1-\frac{1}{c^{\prime}\sqrt{\kappa_{N}}}.$ (46)
@@
In order to ease the notation, we only prove the result for $t=1$. The result for general $t$ follows by linear scaling. We follow the method of proof for estimating the survival probability of a critical Galton–Watson process designed in [44]. Recall the construction of the spinal measure $\mathbb{Q}^{*}$ in Section 5.1. Let us denote by $A_{N}$ the event that the spine is the left-most individual at generation $N$. Recalling that the spine is uniformly distributed among individuals at generation $N$ we have
*$\mathbb{Q}^{*}(Y\mathbbm{1}_{A_{N}})=\mathbb{Q}\Big{(}\frac{Y}{Z_{N}}\Big{)}=\frac{1}{\mu_{N}}\mathbb{P}\left(Y\mathbbm{1}_{Z_{N}>0}\right),$
@@
*$\mathbb{Q}^{*}(Z_{N}\mid A_{N})=\frac{\mathbb{Q}^{*}(Z_{N}\mathbbm{1}_{A_{N}})}{\mathbb{Q}(1/Z_{N})}=\frac{\mu_{N}}{\mathbb{P}\left(Z_{N}>0\right)}.$
We are now left with estimating the expectation on the left-hand side. We need to introduce some notation. Let us denote by $R_{N,i}$ (resp. $L_{N,i}$) the number of individuals at generation $N$ whose most-recent common ancestor with the spine lives at generation $i-1$, and that are to the right (resp. left) of the spine. Let $Z_{N,i}=L_{N,i}+R_{N,i}$ be the total number of individuals that are descended from the offspring of the spine at generation $i$. We have that
@@
and the $(A_{N,i};\,i)$ are independent events. Let us also denote by $R_{i}$ the total number of individuals living at generation $i$ born to the right of the spine, and $\xi_{i}^{*}$ the total number of children of the spine at generation $i$ (including the spine individual at generation $i$). Let $v_{i}$ be the index of the spine at generation $i$. As a preliminary step, we notice that $R_{N,i}$ and $A_{N,i}$ are negatively correlated. Let $Z^{(N-i)}_{j}$ for each $j\geq 1$ be the size at generation $j$ of a BPVE started from a single individual in generation $i$ and stopped at generation $N$. We get
@@
Clearly, $R^{*}_{N}\coloneqq 1+\sum_{i=1}^{N}R^{*}_{N,i}$ has the distribution of $Z_{N}$ under *$\mathbb{Q}^{*}(\cdot\mid A_{N})$. The result will follow by first computing the expectation of $R_{N}\coloneqq 1+\sum_{i=1}^{N}R_{N,i}$ under $\mathbb{Q}^{*}$, then proving that the $L^{1}$ distance of $R_{N}$ and $R^{*}_{N}$ goes to 00 as $N\to\infty$. Let us start by computing the expectation of $R_{N}$. The expected number of children of the spine individual at generation $i-1$ that are not the spine individual at generation $i$ is $f^{\prime\prime}_{i}(1)/f^{\prime}_{i}(1)$. Each of these children ends up to the right of the spine with probability $1/2$, and grows an independent tree with expected size $\mu_{N}/\mu_{i}$. Therefore,
@@
as $N\to\infty$, where the convergence follows from Lemma A.4. Now, let us compare $R_{N}$ to $R^{*}_{N}$. We first have
*$\mathbb{Q}^{*}\Big{[}\frac{1}{\kappa_{N}}|R_{N}-R^{*}_{N}|\Big{]}$ $\leq\frac{1}{\kappa_{N}}\sum_{i=1}^{N}\mathbb{Q}^{*}[R_{N,i}\mathbbm{1}_{\bar{A}_{N,i}}]+\mathbb{Q}^{*}[R^{\prime}_{N,i}]\mathbb{Q}^{*}(\bar{A}_{N,i}).$
Recall that *$\mathbb{Q}^{*}[R^{\prime}_{N,i}]=\mathbb{Q}^{*}[R_{N,i}\mid A_{N,i}]$. Since $R_{N,i}$ and $A_{N,i}$ are negatively correlated (see (48)), we have *$\mathbb{Q}^{*}[R_{N_{i}}\mid A_{N,i}]\leq\mathbb{Q}^{*}[R_{N_{i}}]$. In turn, $R_{N,i}$ and $\bar{A}_{N,i}$ are positively correlated so that
*$\mathbb{Q}^{*}[R^{\prime}_{N,i}]\mathbb{Q}^{*}(\bar{A}_{N,i})=\mathbb{Q}^{*}[R_{N,i}]\mathbb{Q}^{*}(\bar{A}_{N,i})\leq\mathbb{Q}^{*}[R_{N,i}\mathbbm{1}_{\bar{A}_{N,i}}].$
@@
*$\mathbb{Q}^{*}\left[R_{N,i}\mathbbm{1}_{\bar{A}_{N,i}}\right]$ =*[*$=\mathbb{Q}^{*}\bigg{[}\mathbb{Q}^{*}(R_{N,i}\mathbbm{1}_{\bar{A}_{N,i}}\mid\xi^{*}_{i},v_{i})\bigg{]}=\mathbb{Q}^{*}\bigg{[}\mathbb{Q}^{*}(R_{N,i}\mid\xi^{*}_{i},v_{i})\mathbb{Q}^{*}({\bar{A}_{N,i}}\mid\xi^{*}_{i},v_{i})\bigg{]}$
@@
$\mathbb{P}\left(Z_{N}>0\right)=\frac{\mu_{N}}{\mathbb{Q}^{*}[Z_{N}\mid A_{N}]}\sim\frac{\mu_{N}}{\kappa_{N}\frac{1}{\kappa_{N}}\mathbb{Q}^{*}[R_{N}]}\sim\frac{2}{\kappa_{N}}\frac{1}{\int_{[0,1]}e^{-X_{s}}\sigma^{2}(ds)}.$ (49)

### v4

@@ 4.3 The reduced process of a CPP
Recall the definition of the reduced process $(L_{s},0<s<t)$ of an ultrametric tree from Section 3.2.1. The process $L_{s}$ counts the number of ancestral lineages at time $s$, starting from the leaves at $s=0$ and going towards the root at $s=t$. We give a characterization of the reduced process associated with the environmental CPP. It is usual to express the reduced process of a branching process as a time-changed pure birth process, also known as a Yule process.
@@
Let $(L_{s},0<s\leq t)$ be the reduced process associated with the environmental CPP $(0,Z_{e}),d_{\nu_{e}},\operatorname{Leb}]$. Then
@@
Consider a Brownian CPP, that is the random ultrametric space with distribution $\operatorname{CPP}_{1}(\tfrac{dx}{x^{2}},1)$, denoted by $(0,Z_{B}),d_{B},\operatorname{Leb}]$, with corresponding reduced process $(L^{\prime}_{s},s\leq 1)$. It is well-known that
@@
where $(Y(u),u\geq 0)$ is a Yule process, see for instance [10, Lemma 5.5] or [47, Theorem 2.3]. By Proposition 4.2, $(0,Z_{B}),F^{-1}(d_{B}),\bar{\rho}_{t}\operatorname{Leb}]$ is distributed as the environmental CPP (30). Since the identity $F^{-1}(d(x,y))<t\iff d(x,y)<F(t)$ holds, the reduced process $(L_{s},0<s\leq t)$ of $(0,Z_{B}),F^{-1}(d_{B}),\bar{\rho}_{t}\operatorname{Leb}]$ is
@@
and since $L_{t-s}=L^{\prime}_{F(t-s)}\overset{(d)}{=}Y(-\log F(t-s))$ the result follows by noting that
$\forall s\leq t,\quad\frac{1}{F(t-s)}=\frac{2\bar{\rho}_{t}}{\int_{[s,t]}e^{X_{t}-X_{u}}\sigma^{2}(du)}.\qed$
@@ 5.2 The many-to-few formula
where $d_{u}$ denotes the out-degree of $u$. An ultrametric tree with $k$ leaves can always be encoded as a sequence of k1k-1 elements of $\{1,\dots,n\}$ giving the depth of the successive coalescence times between the leaves. More precisely, let $\ell_{1},\dots,\ell_{k}$ be the leaves of $\tau$, ordered such that
@@
where << is the planar (lexicographical) order of the tree. Then $\tau$ can be encoded as the sequence
@@
where $u\wedge v$ denotes the most-recent common ancestor of $u$ and $v$. Conversely, given a sequence $(h_{1},\dots,h_{k-1})\in\{1,\dots,n\}$ one can find an ultrametric tree $\tau$ such that $\mathcal{H}(\tau)=(h_{1},\dots,h_{k-1})$, that is, $\mathcal{H}$ is a bijection. The tree $\mathcal{H}^{-1}(h_{1},\dots,h_{k-1})$ is obtained through the discrete CPP construction illustrated in Figure 3.
@@
Note that there are at most k1k-1 such branch points, and that we have the identity
@@
We now construct the $k$-spine tree using i.i.d. coalescence depths in the CPP construction. More precisely, for each $N$ consider a probability distribution $(p_{i},i\leq tN)$ on $\{1,\dots,tN\}$. (As in the rest of the paper we make the dependency on $N$ implicit.) Let $(H^{(N)}_{i};\,i\geq 1)$ be a sequence of i.i.d. random variables with distribution $(p_{i},i\leq tN)$. The $k$-spine tree $S_{k}$ is defined as the random ultrametric tree $S_{k}=\mathcal{H}^{-1}(H^{(N)}_{1},\dots,H^{(N)}_{k-1})$. As in (19) introduce
@@
Following the notation in [15] we will denote the distribution of this array by $\mathbb{Q}^{k,N}$. Although $\mathbb{Q}^{k,N}$ is formally defined as a measure on $\{0,\dots,tN\}^{k\times k}$, by construction it corresponds to the matrix of pairwise distances between the leaves of the tree $S_{k}$. By abuse of notation, when convenient, we also think of $\mathbb{Q}^{k,N}$ as the law of $S_{k}$, that is as a measure on ultrametric trees.
@@
The reader might be confused by the fact that $\mathbb{Q}$ denotes the law of the size-biased BPVE tree, whereas $\mathbb{Q}^{k,N}$ is the law of a tree $S_{k}$ with a fixed number $k$ of leaves. One could derive an extension of Lemma 5.1 that connects the law $\mathbb{P}$ (biased by its $k$-th factorial moment) to that of a tree with $k$ distinguished lineages obtained by grafting independent subtrees on the spine tree $S_{k}$, see [15, Theorem 3]. In [15] the distribution of this larger tree is denoted by $\bar{\mathbb{Q}}^{k,N}$, and $\mathbb{Q}^{k,N}$ corresponds to the law of the tree spanned by the $k$ distinguished lineages. We follow this notation for easier comparison with [15], even though we do not need such a precise result here.
@@
where $f^{(d_{u})}_{|u|+1}$ denotes the $d_{u}$-th derivative of the generating function of the offspring distribution in generation $|u|$. Note that we have included the $k!$ term in the definition of $\Delta_{k}$ compared to [15]. (As in the rest of the paper, we make the dependency of $\Delta_{k}$ and $(p_{i},i\leq tN)$ on $N$ implicit, but keep it explicit for $\mathbb{Q}^{k,N}$ and $H^{(N)}_{i}$ for later purpose.)
@@ 6 Proof of the Kolmogorov estimate
We aim to prove that $q_{k}\to 1$ as $N\to\infty$, with a uniform control over the speed of convergence for $k\leq(1-\varepsilon)Nt$. This is equivalent to showing that the r.h.s. of (42) diverges uniformly. We have the estimate
@@
by Lemma 1 in [32]. The r.h.s. above can be bounded from below via
$\sum_{j=(1-\varepsilon)Nt}^{Nt-1}\varphi_{j+1}(0)\geq c\sqrt{\kappa_{N}}(\sigma^{2}(t)-\sigma^{2}(t-\varepsilon)),$
by Lemma A.5. Therefore, recalling that by (6) $\mu_{j}$ is uniformly bounded, it follows that, for all $k\leq(1-\varepsilon)Nt$,
$\frac{1}{1-q_{k}}=\frac{1}{\mathbb{P}\left(Z_{Nt}>0\mid Z_{k}=1\right)}\geq\sum_{j=(1-\varepsilon)Nt}^{Nt-1}\frac{\mu_{k}\varphi_{j}(0)}{\mu_{j}}\geq c^{\prime}\sqrt{\kappa_{N}}$
@@
$q_{k}\geq 1-\frac{1}{c^{\prime}\sqrt{\kappa_{N}}}.\qed$
@@
In order to ease the notation, we only prove the result for $t=1$. The result for general $t$ follows by linear scaling. We follow the method of proof for estimating the survival probability of a critical Galton–Watson process designed in [44]. Recall the construction of the spinal measure $\mathbb{Q}^{*}$ in Section 5.1. Let us denote by $A_{N}$ the event that the spine is the left-most individual at generation $N$. Recalling that the spine is uniformly distributed among individuals at generation $N$,
$\mathbb{Q}^{*}(Y\mathbbm{1}_{A_{N}})=\mathbb{Q}\Big(\frac{Y}{Z_{N}}\Big)=\frac{1}{\mu_{N}}\mathbb{P}\left(Y\mathbbm{1}_{Z_{N}>0}\right)$
@@
$\mathbb{Q}^{*}(Z_{N}\mid A_{N})=\frac{\mathbb{Q}^{*}(Z_{N}\mathbbm{1}_{A_{N}})}{\mathbb{Q}(1/Z_{N})}=\frac{\mu_{N}}{\mathbb{P}\left(Z_{N}>0\right)}.$
We are now left with estimating the expectation on the left-hand side.
We need to introduce some notation. Let us denote by $R_{N,i}$ (resp. $L_{N,i}$) the number of individuals at generation $N$ whose most-recent common ancestor with the spine lives at generation i1i-1, and that are to the right (resp. left) of the spine. Let $Z_{N,i}=L_{N,i}+R_{N,i}$ be the total number of individuals that are descended from the offspring of the spine at generation $i$. We have that
@@
and the $(A_{N,i};\,i)$ are independent events. Let us also denote by $R_{i}$ the total number of individuals living at generation $i$ born to the right of the spine, and $\xi_{i}^{*}$ the total number of children of the spine at generation $i$ (including the spine individual at generation $i$). Let $v_{i}$ be the index of the spine at generation $i$.
As a preliminary step, we notice that $R_{N,i}$ and $A_{N,i}$ are negatively correlated. To see this, let $Z^{(N-i)}_{j}$ for each $j\geq 1$ be the size at generation $j$ of a BPVE started from a single individual in generation $i$ and stopped at generation $N$. We get
@@
Clearly, $R^{*}_{N}\coloneqq 1+\sum_{i=1}^{N}R^{*}_{N,i}$ has the distribution of $Z_{N}$ under $\mathbb{Q}^{*}(\cdot\mid A_{N})$. The result will follow by first computing the expectation of $R_{N}\coloneqq 1+\sum_{i=1}^{N}R_{N,i}$ under $\mathbb{Q}^{*}$, then proving that the $L^{1}$ distance of $R_{N}$ and $R^{*}_{N}$ goes to 0 as $N\to\infty$.
Let us start by computing the expectation of $R_{N}$. The expected number of children of the spine individual at generation i1i-1 that are not the spine individual at generation $i$ is $f^{\prime\prime}_{i}(1)/f^{\prime}_{i}(1)$. Each of these children ends up to the right of the spine with probability $1/2$, and grows an independent tree with expected size $\mu_{N}/\mu_{i}$. Therefore,
@@
as $N\to\infty$, where the convergence follows from Lemma A.4.
Now, let us compare $R_{N}$ to $R^{*}_{N}$. We first have
$\mathbb{Q}^{*}\Big[\frac{1}{\kappa_{N}}|R_{N}-R^{*}_{N}|\Big]$ $\leq\frac{1}{\kappa_{N}}\sum_{i=1}^{N}\mathbb{Q}^{*}[R_{N,i}\mathbbm{1}_{\bar{A}_{N,i}}]+\mathbb{Q}^{*}[R^{\prime}_{N,i}]\mathbb{Q}^{*}(\bar{A}_{N,i}).$
Recall that $\mathbb{Q}^{*}[R^{\prime}_{N,i}]=\mathbb{Q}^{*}[R_{N,i}\mid A_{N,i}]$. Since $R_{N,i}$ and $A_{N,i}$ are negatively correlated (see (46)), we have $\mathbb{Q}^{*}[R_{N_{i}}\mid A_{N,i}]\leq\mathbb{Q}^{*}[R_{N_{i}}]$. In turn, $R_{N,i}$ and $\bar{A}_{N,i}$ are positively correlated so that
$\mathbb{Q}^{*}[R^{\prime}_{N,i}]\mathbb{Q}^{*}(\bar{A}_{N,i})=\mathbb{Q}^{*}[R_{N,i}]\mathbb{Q}^{*}(\bar{A}_{N,i})\leq\mathbb{Q}^{*}[R_{N,i}\mathbbm{1}_{\bar{A}_{N,i}}].$
@@
$\mathbb{Q}^{*}\left[R_{N,i}\mathbbm{1}_{\bar{A}_{N,i}}\right]$ $=\mathbb{Q}^{*}\bigg[\mathbb{Q}^{*}(R_{N,i}\mathbbm{1}_{\bar{A}_{N,i}}\mid\xi^{*}_{i},v_{i})\bigg]=\mathbb{Q}^{*}\bigg[\mathbb{Q}^{*}(R_{N,i}\mid\xi^{*}_{i},v_{i})\mathbb{Q}^{*}({\bar{A}_{N,i}}\mid\xi^{*}_{i},v_{i})\bigg]$
@@
$\mathbb{P}\left(Z_{N}>0\right)=\frac{\mu_{N}}{\mathbb{Q}^{*}[Z_{N}\mid A_{N}]}\sim\frac{\mu_{N}}{\kappa_{N}\frac{1}{\kappa_{N}}\mathbb{Q}^{*}[R_{N}]}\sim\frac{2}{\kappa_{N}}\frac{1}{\int_{[0,1]}e^{-X_{s}}\sigma^{2}(ds)}.\qed$